# DCoM: Active Learning for All Learners

### v2

@@ 4.3 Ablation study
We present a series of experiments where we investigate the choices made by our method, including different feature spaces and different uncertainty-based measures. The results demonstrate the robustness of our method to these choices. Additional choices are investigate in App. D, including the contribution of a dynamic $\delta$ and initialization $\delta_{0}$ (App. D.1). In App. D.3, we evaluate the effectiveness of ${S_{\mathcal{L}}}({\mathbb{L}},\Delta)$ from (6) in capturing competence.
Different feature spaces. As discussed in Section 3.1, our approach relies on the existence of a suitable embedding space, where distance is indicative of semantic dissimilarity. We now repeat the basic fully-supervised experiments while varying the embedding employed by DCoM, considering various popular self-supervised representations. Results are reported in Fig. 4, showing that DCoM consistently delivers the best performance regardless of the specific embedding used.
Breaking down the algorithm into its components. We evaluate the contribution of each component in objective function (1) separately. Results are shown in Fig. 4, wherein DPC denotes the Dynamic Probability Coverage, corresponding to ${{\mathcal{O}_{low}}}(x)$ in DCoM, and Margin denotes DCoM’s choice for ${{\mathcal{O}_{high}}}(x)$. Clearly, the weighted objective function of DCoM yields superior results. Similar behavior is observed for several other datasets and other selections of ${{\mathcal{O}_{high}}}(x)$, see App. D.4.
@@
Figure 3: Performance over CIFAR-100, using two different embedding spaces for DCoM and ProbCover, see details in the caption of Fig. 1. Clearly, DCoM consistently achieves the best results. Results using another embedding space are provided in App. D.5.
(c) CIFAR-10
Figure 4: Performance evaluation (mean accuracy and STE) when optimizing each component of objective function (1) separately and together.
Different uncertainty-based methods. We now evaluate our choice for ${{{\mathcal{O}_{high}}}(x)}$ in objective function (1), investigating the alternative methods of min margin, entropy of the min max activation, and uncertainty (see details in Section 4.1). Additionally, we use a score reminiscent of the AL method BADGE [11] and termed Gradient norm, which computes the norm of the gradient between the learner’s last two layers. Once again, the results of DCoM show robustness to this choice, with insignificant differences between the different uncertainty scores, see Fig. 5. Consequently, we prioritize computational efficiency and opt for the fastest-to-compute score, which is min-margin.
Table 3: Model accuracy for different sizes of labeled set ${\mathbb{L}}$ and several uncertainty-based methods for ${\mathbb{F}}(x)$
$|{\mathbb{L}}|$ Random Entropy Margin Uncertainty Gradient norm
200 6.39±0.19 10.71±0.30 10.59±0.19 10.44±0.26 10.33±0.21
400 8.74±0.07 14.35±0.24 14.11±0.11 14.72±0.20 14.59±0.19
600 10.73±0.13 17.07±0.19 16.66±0.11 16.75±0.34 16.83±0.17
1,000 13.49±0.18 19.98±0.22 20.21±0.07 20.40±0.21 20.03±0.17
3,000 24.02±0.27 29.89±0.11 30.36±0.08 30.30±0.11 30.62±0.52
5,000 31.52±0.39 36.08±0.13 36.95±0.10 36.06±0.18 36.12±0.46
7,000 37.69±0.16 41.18±0.25 41.52±0.25 41.80±0.08 41.23±0.65
11,000 45.66±0.22 47.23±0.09 46.86±0.28 46.99±0.27 46.84±0.28
15,000 50.57±0.09 52.67±0.15 52.33±0.22 52.43±0.14 52.21±0.22
19,000 54.95±0.09 56.64±0.27 56.53±0.07 56.47±0.06 56.35±0.28
23,000 57.74±0.14 59.50±0.18 59.60±0.33 59.68±0.24 59.45±0.06
Figure 5: Similar to Fig. 1 and Tables 1-2, using only DCoM while leveraging 4 different variants for its underlying uncertainty score (see text). The performance differences between the variants are small.
@@ Appendix A Implementation details
The source code used in this study is available at https://github.com/avihu111/TypiClust.
The code is based on the GitHub repositories of Yehuda et al. [39] and Munjal et al. [24].
DCoM necessitates the computation of the initial delta for each dataset representation. The values utilized in our experiments can be found in Table 4. The $\delta$ values of CIFAR-10 and CIFAR-100 are sourced from [39]. Information on the computation of $\delta$ values for other datasets can be located in App. D.1.
Table 4: Initial delta values used in experiments
Dataset SSL method Initial $\delta$
@@
Additionally, throughout our experiments we used identical algorithm parameters, distinguishing between those for datasets with smaller class amounts (10) and larger ones (50+). These parameters encompass the adaptive purity threshold $\tau$ from Alg 2 and the logistic function parameters for the competence score ${S_{\mathcal{L}}}({\mathbb{L}},\Delta)$ in (6), $(a,k)$. Specifically, we employed $\tau=0.2~{}\text{cover}+0.4$, with logistic parameters $a=0.9$ and $k=30$ for CIFAR-10, SVHN, and STL-10, and $a=0.8$ and $k=30$ for CIFAR-100 and ImageNet subsets. The $\delta$ resolution for the binary search in DCoM was set to $0.05$. Our ablation study, detailed in App. B, demonstrates that these selections have negligible impact on performance.
@@ A.1 Supervised training
When training on STL-10, SVHN, CIFAR-10 and CIFAR-100 datasets, we utilized a ResNet-18 architecture trained for 200 epochs. Our optimization strategy involved using an SGD optimizer with a Nesterov momentum of 0.9, weight decay set to 0.0003, and cosine learning rate scheduling starting at a base rate of 0.025. Training was performed with a batch size of 100 examples, and we applied random cropping and horizontal flips for data augmentation. For an illustration of these parameters in use, refer to [24].
@@ A.2 Semi-supervised training
When training FlexMatch [42], we used the semi supervised framework by [35], and the AL framework by [24]. All experiments involved 3 repetitions.
We relied on the standard hyper-parameters used by FlexMatch [42]. Specifically, we trained WideResNet-28 for 512 epochs using the SGD optimizer, with $0.03$ learning rate, $64$ batch size, $0.9$ SGD momentum, $0.999$ EMA momentum, $0.001$ weight decay and $2$ widen factor.
@@ A.3 Self-supervised feature extraction
STL-10, SVHN, CIFAR-10, CIFAR-100. To extract semantically meaningful features, we trained SimCLR using the code provided by [33] for STL-10, SVHN, CIFAR-10 and CIFAR-100. Specifically, we used ResNet-18 [16] with an MLP projection layer to a $128$-dim vector, trained for 512 epochs. All the training hyper-parameters were identical to those used by SCAN (all details can be found in [33]). After training, we used the $512$ dimensional penultimate layer as the representation space.
@@
Representation learning: ImageNet-50/100/200. We extracted features from the official (ViT-S/14) DINOv2 weights pre-trained on ImageNet [26]. We employed the L2-normalized penultimate layer for the embedding, which has a dimensionality of $384$.
@@
Ablation embeddings – CIFAR-100. We extracted more features as BYOL [13], MOCOv2+ [17], and Barlow Twins [41] using pre-trained weights from [8].
Appendix B Hyper-parameters exploring
@@
As described before, DCoM has 3 hyper-parameters: $a,k\text{and }\tau$. When running the main experiments described in Section 4.2, we maintained consistency by employing identical parameters for STL-10, SVHN, and CIFAR-10, as well as for CIFAR-100 and the subsets of ImageNet. This approach helps to demonstrate that the choice of the hyper-parameters is not critical and does not significantly affect the results. Here, we conduct experiments with DCoM using STL-10 and CIFAR-100 datasets and several values for $a$ and $k$ from the definition of the competence score ${S_{\mathcal{L}}}({\mathbb{L}},\Delta)$ in (6). We observe that while certain choices may be slightly better than others, all choices yield results that are very close to each other, and DCoM consistently outperforms or matches the performance of previous methods.
@@
Figure 6: Evaluation of DCoM across different settings of the logistic function hyper-parameters a and k. Each plot displays the mean and standard error over $3-5$ repetitions. The findings indicate that fine-tuning these parameters has negligible effects on performance.
Appendix C Time and space complexity of DCoM
@@
For the complexity analysis below, let $n$ represent the number of examples in the combined unlabeled and labeled pool $|{\mathbb{L}}\cup{\mathbb{U}}|$, $d$ the dimension of the data embedding space, $b$ the given budget ($|{\mathbb{L}}|$) and $q$ the size of the query set. As previously stated, DCoM can be divided into three distinct steps:
C.1 Selection of initial $\boldsymbol{\delta}$
@@
Time Complexity: The process begins by creating $t$ adjacency graphs, each corresponding to a different $\delta$ value being evaluated. Here, $t$ denotes the number of $\delta$’s being examined. The time complexity of each graph generation is $O(n^{2}\cdot d)$ time. The computation of purity requires the prediction of pseudo-labels by k-means at a cost of $O(n^{2})$. The complexity of computing ball purity for each example is $O(n^{2})$, so totally each the ball purity computing takes $O(n^{3})$. This adds up to overall time complexity of $O(tdn^{3})$. Although it may seem significant, this step doesn’t occur during the active learning algorithm process. Additionally, it’s possible to confine the data employed for purity computation to the densest points, effectively lowering the complexity to $O(tdn^{2})$. This approach mandates the use of at least the number of points equivalent to the number of classes for purity computation. This technique yields similar $\delta_{0}$ values (consistent for CIFAR-10, CIFAR-100, and ImageNet-50, and with a variation of 0.05 for ImageNet-100/200 and STL-10 and 0.15 for SVHN).
@@
Space complexity: Naively, the space complexity is $O(n^{2})$, which might be impractical for large datasets like ImageNet. However, [39] demonstrates that utilizing a sparse matrix in coordinate list (COO) format and using limited $\delta$ values, the space complexity becomes $O(|E|)$, where $E$ represents the set of edges in the graph. Although $O(|E|)$ remains $O(n^{2})$ in the worst-case scenario, in practice, the average vertex degree with a limited radius is less than $n$.
C.2 Customizing $\boldsymbol{\delta}$ values
@@
Time complexity: Each iteration of DCoM starts with customizing $\delta$ values for each labeled example from the last active step. In the worst case, there are $|{\mathbb{L}}|$ examples. Customizing $\delta$ involves building the adjacency graph using $\delta_{\text{max}}$, running all examples on the model to obtain their predictions, and applying binary search over $\delta$ values between 0 and $\delta_{\text{max}}$ for each example from the last active step. The binary search is performed on a continuous parameter ($\delta$) with a search resolution $r$. This implies that there are $\frac{\delta_{\text{max}}}{r}$ available values for $\delta$. Overall, this process takes $O(dn^{2}+n+|{\mathbb{L}}|\cdot\log(\frac{\delta_{\text{max}}}{r}))=O(dn^{2}+b\cdot\log(\frac{\delta_{\text{max}}}{r}))$. In practical terms, due to the vectorization of these processes, it requires approximately 27 minutes to construct an adjacency graph for ImageNet-200 and update deltas for 1000 examples on a single CPU.
Space complexity: Similar to the space complexity analysis in App. C.1, which is $O(n^{2})$.
C.3 Active sampling

### v3

@@ 4.3 Ablation study
We present a series of experiments where we investigate the choices made by our method, including different feature spaces and different uncertainty-based measures. The results demonstrate the robustness of our method to these choices. Additional choices are investigate in Appendix E, including the contribution of a dynamic $\delta$ and initialization $\delta_{0}$ (Appendix E.1). In Appendix E.3, we evaluate the effectiveness of ${S_{\mathcal{L}}}({\mathbb{L}},\Delta)$ from (5) in capturing competence.
@@
Figure 5: Performance over CIFAR-100, using two different embedding spaces for DCoM and ProbCover, see details in the caption of Fig. 3. Clearly, DCoM consistently achieves the best results. Results using an additional embedding space are shown in Appendix E.4.
Different feature spaces. As discussed in Section 3, our approach relies on the existence of a suitable embedding space, where distance is indicative of semantic dissimilarity. We now repeat the basic fully-supervised experiments while varying the embedding space, comparing various popular self-supervised representations. Results are reported in Fig. 5, showing that DCoM consistently delivers the best performance regardless of the specific embedding used.
Breaking down the algorithm into its components. The objective function in (1) includes the linear combination of two objective functions, weighed by competence score ${S_{\mathcal{L}}}({\mathbb{L}},\Delta)$ from (5). We evaluate the contribution of each component in this objective function separately. Results are shown in Fig. 6, wherein DPC denotes the Dynamic Probability Coverage, corresponding to ${{\mathcal{O}_{low}}}(x)$ in DCoM, and Margin denotes DCoM’s choice for ${{\mathcal{O}_{high}}}(x)$. Clearly, the weighted objective function of DCoM yields superior results.
(a) CIFAR-10
(b) IMAGENET-50
Figure 6: Performance (mean accuracy and STE) when optimizing each component of objective function (1) separately and together.
Different uncertainty-based methods. We evaluate our choice for ${{\mathcal{O}_{high}}}(x)$} in objective function (1), investigating the alternative methods of min margin, entropy of the min max activation, and uncertainty (see details in Section 4.1). Additionally, we use a score reminiscent of the AL method BADGE (Gal, Islam, and Ghahramani 2017) and termed Gradient norm, which computes the norm of the gradient between the learner’s last two layers. Once again, the results of DCoM show robustness to this choice, with insignificant differences between the different uncertainty scores, see Fig. 7. Consequently, we prioritize computational efficiency and opt for the fastest-to-compute score, which is min-margin.
Figure 7: Similar to Fig. 3, using only DCoM while leveraging 4 different variants for its underlying uncertainty score (see text). The performance differences between the variants are hardly significant.
@@ 5 Summary
Acknowledgments
This work was supported by grants from the Israeli Council of Higher Education, AFOSR award FA8655-24-1-7006, and the Gatsby Charitable Foundation.
@@ Appendix A Implementation details
The source code for this study can be found in the zip file included in the supplementary material. This zip file contains a USAGE.md file that provides instructions on how to run the algorithm. The code in the zip file is based on the GitHub repositories of Yehuda et al. (2022) and Munjal et al. (2022). Our GitHub repository will be made public upon acceptance.
DCoM necessitates the computation of the initial delta for each dataset representation. The values utilized in our experiments can be found in Table 2. The $\delta$ values of CIFAR-10 and CIFAR-100 are sourced from (Yehuda et al. 2022). Information on the computation of $\delta$ values for other datasets can be located in Appendix E.1.
Dataset SSL method $\delta_{0}$
@@
Table 2: Initial delta values used in experiments
Additionally, throughout our experiments we used identical algorithm parameters, distinguishing between those for datasets with smaller class amounts (10) and larger ones (50+). These parameters encompass the adaptive purity threshold $\tau$ from Alg 2 and the logistic function parameters for the competence score ${S_{\mathcal{L}}}({\mathbb{L}},\Delta)$ in (5), $(a,k)$. Specifically, we employed $\tau=0.2~\text{cover}+0.4$, with logistic parameters $a=0.9$ and $k=30$ for CIFAR-10, SVHN, and STL-10, and $a=0.8$ and $k=30$ for CIFAR-100 and ImageNet subsets. The $\delta$ resolution for the binary search in DCoM was set to $0.05$. Our ablation study, detailed in Appendix C, demonstrates that these selections have negligible impact on performance.
@@ A.1 Supervised training
When training on STL-10, SVHN, CIFAR-10 and CIFAR-100 datasets, we utilized a ResNet-18 architecture trained for 200 epochs. Our optimization strategy involved using an SGD optimizer with a Nesterov momentum of 0.9, weight decay set to 0.0003, and cosine learning rate scheduling starting at a base rate of 0.025. Training was performed with a batch size of 100 examples, and we applied random cropping and horizontal flips for data augmentation. For an illustration of these parameters in use, refer to (Munjal et al. 2022).
@@ A.2 Semi-supervised training
When training FlexMatch (Zhang et al. 2021), we used the semi supervised framework by (Wang et al. 2022), and the AL framework by (Munjal et al. 2022). All experiments involved 3 repetitions.
We relied on the standard hyper-parameters used by FlexMatch (Zhang et al. 2021). Specifically, we trained WideResNet-28 for 512 epochs using the SGD optimizer, with $0.03$ learning rate, $64$ batch size, $0.9$ SGD momentum, $0.999$ EMA momentum, $0.001$ weight decay and $2$ widen factor.
@@ A.3 Self-supervised feature extraction
STL-10, SVHN, CIFAR-10, CIFAR-100. To extract semantically meaningful features, we trained SimCLR using the code provided by (Van Gansbeke et al. 2020) for STL-10, SVHN, CIFAR-10 and CIFAR-100. Specifically, we used ResNet-18 (He et al. 2016) with an MLP projection layer to a $128$-dim vector, trained for 512 epochs. All the training hyper-parameters were identical to those used by SCAN (all details can be found in (Van Gansbeke et al. 2020)). After training, we used the $512$ dimensional penultimate layer as the representation space.
@@
Representation learning: ImageNet-50/100/200. We extracted features from the official (ViT-S/14) DINOv2 weights pre-trained on ImageNet (Oquab et al. 2023). We employed the L2-normalized penultimate layer for the embedding, which has a dimensionality of $384$.
@@
Ablation embeddings – CIFAR-100. We extracted more features as BYOL (Grill et al. 2020), MOCOv2+ (He et al. 2020), and Barlow Twins (Zbontar et al. 2021) using pre-trained weights from (da Costa et al. 2022).
Appendix B Purity Calculation in nNN
When transitioning from 1NN to nNN, the structure of the problem remains unchanged, with the key difference being that the radius $\delta$ is adjusted for each labeled example. Next, we demonstrate that the purity of each ball in the nNN problem can be precisely computed using these $\delta$-values. This is because when a labeled example $c\in{\mathbb{L}}$ assigns a label to a point $x$, it must also cover $x$. Therefore, any mislabeling by $c$ would reduce the purity, as illustrated by the following statement:
Proposition.
If $c\in{\mathbb{L}}$ is labeling $x$ and $x\in C({\mathbb{L}},\Delta)$, then $x\in B_{\delta_{c}}(c)$.
Proof.
Let $c\in{\mathbb{L}}$ be the nearest neighbor (nNN) of $x$, and let $\delta_{c}$ be the corresponding radius for $c$. By definition, $f^{nNN}(x)=f(x)$, and since $c$ is labeled, we have $f(c)=f^{nNN}(c)\equiv f^{N}(c)$.
Given that $x\in C({\mathbb{L}},\Delta)$, there exists a labeled example $c^{\prime}\in{\mathbb{L}}$ that covers $x$, meaning $d(x,c^{\prime})<\delta_{c^{\prime}}$. Suppose that $c\neq c^{\prime}$. Since $c$ is the nNN of $x$, it follows that:
$\frac{d(x,c)}{\delta_{c}}<\frac{d(x,c^{\prime})}{\delta_{c^{\prime}}}$
Given $d(x,c^{\prime})<\delta_{c^{\prime}}$, we get:
$\frac{d(x,c)}{\delta_{c}}<1\Rightarrow d(x,c)<\delta_{c}$
Thus, $x\in B_{\delta_{c}}(c)$, meaning $c$ covers $x$ within the radius $\delta_{c}$.
This implies that if $x$ were wrongly labeled, $c$ would still cover $x$, indicating lower purity in $B_{\delta_{c}}(c)$. ∎
Appendix C Hyper-parameters exploring
Theoretical basis for the Choice of $\boldsymbol{a}$ and $\boldsymbol{k}$. The parameter $a$ represents the center of the sigmoid function. Our objective is for the algorithm to assign increased weight to $\mathcal{O}_{high}$} only after it has covered a substantial portion of the dataset. Consequently, we choose $a\in(0.5,1)$. The parameter $k$ controls the steepness of the sigmoid curve. The function must yield 0 in the absence of coverage and approach 1 as coverage becomes complete. To achieve this, we configure the numerator of the fraction to be equal to the denominator when the cover’s probability is 1. Although the function cannot pass precisely through $(0,0)$, it can asymptotically approach $(0,\varepsilon)$ for an infinitesimally small $\varepsilon\in\Re$.
We now investigate the relationship between this approximation error and the parameter $k$. Define the function $f(x)$ as $S_{\mathcal{L}}$} for simplicity:
$f(x)=\frac{1+e^{-k(1-a)}}{1+e^{-k(x-a)}}$
The limit of $f(x)$ as $x$ approaches 0 is
$\lim_{x\to 0}f(x)=\frac{1+e^{-k(1-a)}}{1+e^{ka}}$
For the function to approach 0, we require:
$\frac{1+e^{-k(1-a)}}{1+e^{ka}}\ll 1$
This implies:
$1+e^{-k(1-a)}\ll 1+e^{ka}$
$e^{-k(1-a)}\ll e^{ka}$
$e^{-k}\ll e^{ka}$
$e^{-k}\ll 1$
Next, we calculate $f(0)$:
$f(0)=\frac{1+e^{-k(1-a)}}{1+e^{ka}}=\varepsilon$
Solving for $k$:
$1+e^{-k(1-a)}=\varepsilon(1+e^{ka})$
$1-\varepsilon=\varepsilon e^{ka}-e^{-k(1-a)}$
$1-\varepsilon=e^{ka}(\varepsilon-e^{-k})$
$e^{ka}=\frac{1-\varepsilon}{\varepsilon-e^{-k}}$
Approximating $e^{ka}$ using (*):
$e^{ka}\approx\frac{1-\varepsilon}{\varepsilon}$
So:
$ka\approx\ln\left(\frac{1-\varepsilon}{\varepsilon}\right)$
Finally:
$k\approx\frac{1}{a}\ln\left(\frac{1}{\varepsilon}-1\right)\approx\frac{\ln\left(\frac{1}{\varepsilon}\right)}{a}$
To minimize the approximation error, we should choose a larger $k$. In our experiments with $a=0.8\text{ or }0.9$ and $\varepsilon=10^{-10}$, we determined:
$k(a=0.9,\varepsilon=10^{-10})\approx 25.58$
$k(a=0.8,\varepsilon=10^{-10})\approx 28.78$
The value of $\varepsilon$ was chosen to be asymptotically smaller than the dimension of the initial graphs constructed for all the datasets. In accordance, we chose $k=30$.
@@ Appendix C Hyper-parameters exploring
Experimental assessment of DCoM hyper-parameter consistency As described before, DCoM has 3 hyper-parameters: $a,k\text{ and }\tau$. When running the main experiments described in Section 4.2, we maintained consistency by employing identical parameters for STL-10, SVHN, and CIFAR-10, as well as for CIFAR-100 and the subsets of ImageNet. This approach helps to demonstrate that the choice of the hyper-parameters is not critical and does not significantly affect the results. Here, we conduct experiments with DCoM using STL-10 and CIFAR-100 datasets and several values for $a$ and $k$ from the definition of the competence score ${S_{\mathcal{L}}}({\mathbb{L}},\Delta)$ in (5). We observe that while certain choices may be slightly better than others, all choices yield similar results, and DCoM consistently outperforms or matches the performance of previous methods.
@@
Figure 8: Evaluation of DCoM across different settings of the logistic function hyper-parameters a and k. Each plot displays the mean and standard error over 353-5 repetitions. The findings indicate that fine-tuning these parameters has negligible effects on performance.
Appendix D Time and space complexity of DCoM
@@
For the complexity analysis below, let $n$ represent the number of examples in the combined unlabeled and labeled pool $|{\mathbb{U}}\cup{\mathbb{L}}|$, $d$ the dimension of the data embedding space, $b$ - the given budget ($|{\mathbb{L}}|$) and $q$ - the size of the query set. As previously stated, DCoM can be divided into three distinct steps:
D.1 Selection of initial $\boldsymbol{\delta}$ delta
@@
Time Complexity: The process begins by creating $t$ adjacency graphs, each corresponding to a different $\delta$ value being evaluated. Here, $t$ denotes the number of $\delta$’s being examined. The time complexity of each graph generation is $O(n^{2}\cdot d)$ time. The computation of purity requires the prediction of pseudo-labels by k-means at a cost of $O(n^{2})$. The complexity of computing ball purity for one example is $O(n^{2})$, so totally the ball purity computing for one graph takes $O(n^{3})$. This adds up to overall time complexity of $O(tdn^{2}+n^{2}+tn^{3})=O(tn^{3})$. We can construct the initial graph using the largest $\delta$-value and then, instead of building a new graph for each $\delta$-value in $O(tdn^{2})$, we can filter distances above the current $\delta$ in $O(dn^{2}+n^{2})$. However, the total complexity remains $O(tn^{3})$. Although it may seem significant, this step doesn’t occur during the active learning algorithm process. Additionally, it’s possible to confine the data employed for purity computation to the densest points, effectively lowering the complexity to $O(dn^{2})$. This approach mandates the use of at least the number of points equivalent to the number of classes for purity computation. This technique yields similar $\delta_{0}$ values (consistent for CIFAR-10, CIFAR-100, and ImageNet-50, and with a variation of 0.05 for ImageNet-100/200 and STL-10 and 0.15 for SVHN).
@@
Space complexity: Naively, the space complexity is $O(n^{2})$, which might be impractical for large datasets like ImageNet. However, (Yehuda et al. 2022) demonstrates that utilizing a sparse matrix in coordinate list (COO) format and using limited $\delta$ values, the space complexity becomes $O(|E|)$, where $E$ represents the set of edges in the graph. Although $O(|E|)$ remains $O(n^{2})$ in the worst-case scenario, in practice, the average vertex degree with a limited radius is less than $n$.
D.2 $\boldsymbol{\Delta}$-Adjustment
@@
Time complexity: Each iteration of DCoM starts with customizing $\delta$ values for each labeled example from the last active step. In the worst case, there are $|{\mathbb{L}}|$ examples. Customizing $\Delta$ involves building the adjacency graph using $\delta_{\text{max}}$, running all examples on the model to obtain their predictions, and applying binary search over $\delta$ values between 0 and $\delta_{\text{max}}$ for each example from the last active step. The binary search is performed on a continuous parameter ($\delta$) with a search resolution $r$. This implies that there are $\frac{\delta_{\text{max}}}{r}$ available values for $\delta$. For each $\delta$ value, the algorithm computes its purity value. Overall, this process takes $O(dn^{2}+n+|{\mathbb{L}}|\cdot\log(\frac{\delta_{\text{max}}}{r})\cdot n^{2})=O(dn^{2}+b\cdot\log(\frac{\delta_{\text{max}}}{r})\cdot n^{2})$. In practical terms, due to the vectorization of these processes, it requires approximately 27 minutes to construct an adjacency graph for ImageNet-200 and update deltas for 1000 examples on a single CPU.
Space complexity: Similar to the space complexity analysis in Appendix D.1, which is $O(n^{2})$.
D.3 Active sampling